OpenStudy (anonymous):

Jacob and Roman are standing next to each other. Jacob, who is 152 cm tall, has a shadow that is 38 cm long. If Roman is 136 cm tall, approximately how long is his shadow? 16 cm 34 cm 42 cm 54 cm

OpenStudy (anonymous):

You can set up a proportion to determine the shadow length: 38/152 = x/136 Now, just cross-multiply and divide: 152x = (38)(136) -> x = (38)(136)/152

OpenStudy (anonymous):

All good now, @stepennd7 ?

OpenStudy (anonymous):

b 34?

OpenStudy (anonymous):

Yes, you got it! Good job!